OpenStudy (anonymous):

The volume of a regular triangular prism is 90 sqrt of 3 cm cube. If the radius of the inscribed circle in the base is 3cm, find the altitude of the prism. Please help. Including the drawing. Thanks.

OpenStudy (sepeario):

What do you mean by "90 sqrt of 3 cm cube"?

OpenStudy (anonymous):

r=3 d=2*3=6 edge of square base=6 \[volume=\frac{ 1 }{ 3 }\pi a^2h\] \[\frac{ 1 }{ 3 }\pi 6^2h=90\sqrt{3}\] h=?

OpenStudy (anonymous):

Cm is raise to cube

OpenStudy (anonymous):

90 sqrt of 3 cm^3